### v1

# AutoBS: Autonomous Base Station Deployment Framework with Reinforcement Learning and Digital Twin Network

Abstract
This paper introduces AutoBS, a reinforcement learning (RL)-based framework for optimal base station (BS) deployment in 6G networks. AutoBS leverages the Proximal Policy Optimization (PPO) algorithm and fast, site-specific pathloss predictions from PMNet to efficiently learn deployment strategies that balance coverage and capacity. Numerical results demonstrate that AutoBS achieves 95% for a single BS, and 90% for multiple BSs, of the capacity provided by exhaustive search methods while reducing inference time from hours to milliseconds, making it highly suitable for real-time applications. AutoBS offers a scalable and automated solution for large-scale 6G networks, addressing the challenges of dynamic environments with minimal computational overhead.
I Introduction
The rollout of 6G networks demands higher base station (BS) density due to the use of higher frequencies like millimeter-wave (mmWave), which offers enhanced bandwidth and low latency. However, these frequencies suffer from severe signal attenuation and limited propagation range, particularly in complex urban environments. As a result, dense BS deployment becomes essential to maintain reliable coverage and capacity. Furthermore, cell-free massive MIMO, where a significant number of access points with one of a few antennas each are distributed over an area [1] is anticipated to be widely used in 6G.
However, optimizing BS placement in such environments presents significant challenges due to highly site-specific conditions. Traditional BS deployment methods often rely on manual planning, based on real-world measurement of the propagation conditions [2], and/or computationally intensive ray-tracing simulations using tools like WirelessInsite [3] or SionnaRT [4]. The fact that these approaches are time- and labor- consuming makes them less suited for the dense deployment in 6G. Furthermore, they are not suited for real-time adaptation of network topology, e.g., dynamic addition and placement of (mobile) BSs in reaction to change of user density, e.g., at special events.
Such network planning can be also seen as a special case of Digital Twin (DT) networks, which create virtual replicas of physical environments, enabling real-time simulation and optimization of network performance under site-specific conditions. Again, manually optimizing BS placement within DT frameworks remains computationally expensive.
Therefore, large-scale and real-time network optimization requires the integration of machine learning (ML)-based approaches to automate and expedite the process.
To address these challenges, we propose AutoBS, an autonomous BS deployment framework that uses Deep Reinforcement Learning (DRL) and a DT generator to optimize BS placement in a site-specific manner. AutoBS integrates PMNet [5, 6], a fast and accurate ML-based path gain predictor, to calculate reward based on both coverage and capacity metrics. By autonomously learning optimal placement strategies, AutoBS significantly reduces the computational overhead and deployment time compared to conventional methods, making it well-suited for real-time and large-scale 6G network optimization.
Contributions. This paper presents AutoBS, an autonomous BS deployment framework utilizing DRL and a DT generator (PMNet). Our key contributions are summarized as follows:
  • AutoBS Framework Design: We introduce a novel DRL-based framework for BS deployment that incorporates PMNet for real-time, site-specific channel predictions [^1] Selected results and codes are available at github.com/abman23/autobs. PMNet enables fast, precise reward calculations, allowing AutoBS to effectively optimize coverage and capacity through an adaptive reward function (see Fig. 2 in Sec. III). Details on reward design are provided in Table V in Sec. VI.
  • Support for Multi-BS Deploy.: AutoBS handles both static single BS and asynchronous multi-BS deployments, providing flexibility across diverse deployment scenarios (see Fig. 5 in Sec. IV).
  • Fast and Near-Optimal Deployment: Simulations demonstrate that AutoBS reduces inference time from hours to milliseconds compared to exhaustive methods, particularly in multi-BS deployments, making it practical for large-scale, real-time optimization (see Table IV in Sec. IV).
II Methodology
II-A Problem Formulation and DRL Approach
The BS deployment problem can be mathematically formulated as an optimization task. Let the BS location be represented by coordinates $(i,j)$ on a site-specific building map $m$. The objective function to optimize is:
$\max_{\{i,j\}}\sum_{m=1}^{M}\left(V_{m}+\nu C_{m}\right),\ \text{s.t.}\ \{i,j\}\in\mathcal{B},$ (1)
where $V_{m}$ is the coverage, $C_{m}$ represents the capacity, and $\nu$ is a weighting factor that governs the trade-off between coverage and capacity. The set $\mathcal{B}$ defines the deployable areas for BS placement, such as rooftops or other designated locations within the environment.
AutoBS frames the deployment problem as a reinforcement learning task, where a DRL agent interacts with the environment to learn optimal placement strategies. Through continuous feedback via rewards, the agent refines its decisions, capturing site-specific channel characteristics and optimizing network performance over time.
Figure 1: Overview of AutoBS framework.
II-B Reward Calculation via PMNet
In DRL, efficient reward calculation is essential for guiding the learning process, where the agent explores numerous actions throughout training. However, calculating network performance metrics, such as coverage and capacity, based on BS placements—key components of the reward function—is computationally intensive. Traditional RT simulations, which can take tens of minutes to generate a single pathloss map, are impractical for real-time reinforcement learning.
To overcome this challenge, AutoBS integrates the PMNet framework, a DT generator, for rapid reward computation. PMNet predicts pathloss maps with an RMSE on the order of $10^{-2}$ within milliseconds, enabling the Proximal Policy Optimization (PPO) algorithm [7] to immediately evaluate network performance and receive reward feedback after each deployment decision. This fast feedback allows the agent to efficiently simulate and assess multi-BS deployment scenarios, significantly accelerating training by exploring more placement options with timely updates. The integration of PMNet ensures both speed and scalability, enabling the AutoBS to generalize effectively across diverse environments.
III AutoBS: An Autonomous Base Station Deployment Framework
III-A Architecture
The AutoBS framework automates optimal BS placement in 6G networks using DRL, with a PPO agent. PPO is well-suited for this task due to its balance between exploration (testing new deployment locations) and exploitation (refining placement strategies). By using a clipped objective function, PPO ensures stable policy updates and prevents large, destabilizing changes during training.
The framework supports two deployment modes: (1) static single BS deployment, where a single BS is placed in a fixed environment, and (2) asynchronous multi-BS deployment, where multiple BSs are deployed asynchronously.
III-A1 Single (Static) BS Deployment
In the static single BS deployment scenario, the geographical environment and network conditions are assumed to be fixed. The input to the system is a site-specific map $\mathcal{S}$, which contains detailed information about buildings, terrain, and obstacles. The PPO agent processes this map and outputs the optimal location $(i,j)$ for placing the BS to maximize coverage and capacity.
III-A2 Multi (Asynchronous) BS Deployment
In the asynchronous multi-BS deployment scenario, BSs are deployed incrementally over time. After each deployment, the environment is updated to reflect the changes in network conditions, such as coverage, capacity, and user distribution. The PPO agent adjusts its strategy after each deployment based on real-time feedback, refining its decisions progressively.
III-B Evaluation Metrics
The AutoBS framework’s performance is evaluated using two key metrics: Coverage and Capacity. These metrics are essential for assessing the quality of BS placements in terms of both spatial coverage and network throughput.
Coverage. represents the proportion of the area or users receiving sufficient signal strength, e.g., received power, from the deployed BSs. It is calculated as the total area where the received power $P_{i,j}$ exceeds a predefined threshold $\mathrm{thr}$, ensuring satisfactory Quality of Service (QoS). The coverage $V$ is expressed as:
$V=\sum_{\{i,j\}\in\mathcal{R}}{v_{i,j}},$ (2)
where $\mathcal{R}$ is the region of interest (RoI), e.g., non-deployable and non-building areas, and $v_{i,j}$ is a binary indicator of coverage:
$v_{i,j}=\begin{cases}1,&P_{i,j}\geq\mathrm{thr},\\ 0,&P_{i,j}<\mathrm{thr}.\end{cases}$ (3)
The threshold $\mathrm{thr}$ depends on network QoS requirements (e.g., the minimum signal strength for reliable communication).
Capacity. represents the maximum amount of data that can be transmitted or received over a network, which depends on the quality of the connection, e.g., signal-to-noise ratio (SNR). The total capacity $C$ is given by:
$C=\sum_{\{i,j\}\in\mathcal{R}}\log_{2}{\left(1+\text{SNR}_{i,j}\right)},$ (4)
where $\text{SNR}_{i,j}$ is the signal-to-noise ratio at pixel $(i,j)$, calculated as:
$\text{SNR}_{i,j}=\frac{\sum_{k}P^{k}_{i,j}}{\sigma^{2}}.$ (5)
Here, $P^{k}_{i,j}$ represents the received power from the $k$-th BS at pixel $(i,j)$, and $\sigma^{2}$ is the noise variance. In our simulations, we chose $\sigma^{2}=\eta/4$, ensuring that the SNR at the coverage boundary is at least 6 dB. [^2] In an actual DT simulation, noise variance would obviously follow from the system bandwidth, transmit power, receiver noise figure, and antenna again; but since cell-edge SNR of 6 dB is a common planning goal, we directly used it here).
III-C Training
The AutoBS framework models the base station (BS) deployment problem as a Markov Decision Process (MDP), enabling the agent to interact with its environment to make optimal BS placement decisions.
III-C1 MDP Design
Environment. The environment for the PPO agent is defined by the interactions within a site-specific building map, where each deployed BS impacts the network performance metrics such as coverage and capacity. At each time step $t$, the agent interacts with this environment by observing a state $s_{t}\in\mathcal{S}$, which includes details on buildings, obstacles, and terrain specific to the deployment site. The agent then takes an action $a_{t}\in\mathcal{A}$, selecting a BS deployment location based on its learned policy $\pi$. After the action is executed, the environment transitions to a new state $s_{t+1}$, representing the updated network layout and performance with the new BS in place. The agent receives a reward $r_{t}$ based on the effectiveness of its decision, guiding it toward an optimal deployment policy $\pi^{*}$.
State. The state $s_{t}$ at time step $t$ provides key information for decision-making about BS placement. Formally, the state is expressed as:
$s_{t}=\left\{\mathcal{S}\right\},$ (6)
where $\mathcal{S}$ represents the site-specific building map. This map encompasses critical details about building locations, obstacles, and terrain, all of which significantly influence signal propagation (i.e., site-specific channel).
Action. The action space $\mathcal{A}$ consists of potential BS deployment locations. At each time step $t$, the agent selects an action $a_{t}\in\mathcal{A}$, which corresponds to placing a BS at a specific geographical coordinate $(i,j)$ on the site-specific map. The action is represented as:
$a_{t}=(i,j),\quad\{i,j\}\in\mathcal{B},$ (7)
where $\mathcal{B}$ is the set of all permissible deployment locations within the map. The agent deploys BSs sequentially, choosing new locations at each time step based on current network needs. This approach is computationally efficient, allowing the agent to adapt its strategy in real-time as it learns from previous decisions.
Reward. The reward function $r_{t}$ incentivizes the agent to improve network performance by maximizing both coverage and capacity. The reward at time step $t$ is defined as:
$r_{t}=\nu_{1}V_{t}+\nu_{2}C_{t}+\nu_{3}P_{t},$ (8)
where $V_{t}$ and $C_{t}$ represent the improvements in coverage and capacity, respectively, and $P_{t}=\sum_{\{i,j\}\in\mathcal{R}}P_{i,j}$ denotes the total pathgain, for the $t$-th deployment. The weighting parameter $\nu$ adjusts the relative importance of coverage, capacity, and pathgain in the overall network optimization. For further details regarding the reward design, please refer to Sec. VI. Note also that for more general DT applications (e.g., incorporating user demands), only the reward function needs to be adapted, while the remaining framework of AutoBS stays the same.
As mentioned in Sec. II-B, the reward must be recalculated after each BS deployment to reflect the updated network conditions. This typically requires coverage and capacity evaluation from pathloss map prediction, which are computationally expensive. To overcome this challenge, the AutoBS framework integrates the PMNet model, which provides fast and accurate pathloss predictions, enabling efficient reward calculations. PMNet allows the agent to quickly assess the network performance after each BS placement, ensuring that the training process is both realistic and computationally feasible.
Figure 2: Training for AutoBS framework.
III-C2 Training Process
The PPO agent is trained using the PPO algorithm [7], which allows the agent to iteratively interact with a simulated environment. The agent selects BS deployment actions and receives rewards based on improvements in network coverage and capacity, progressively refining its policy $\pi_{\theta}(a_{t}|s_{t})$ over time.
At each time step $t$, the agent updates its policy to maximize the cumulative reward:
$J(\theta)=\mathbb{E}\left[\sum_{t=0}^{T}\gamma^{t}r_{t}\right],$ (9)
where $\gamma$ is the discount factor, $r_{t}$ is the reward received at time step $t$, and $T$ is the time horizon over which the agent aims to optimize its decisions. The discount factor $\gamma$ balances immediate and future rewards, promoting long-term planning in BS placement.
The PPO agent optimizes the following objective function:
$\mathcal{L}(\theta)=\mathbb{E}_{t}\left[\min\left(r_{t}(\theta)\hat{A}_{t},\,\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}\right)\right],$ (10)
where $r_{t}(\theta)$ is the probability ratio between the updated and previous policies, $\epsilon$ is a clipping parameter to limit policy changes for stability, and $\hat{A}_{t}$ represents the advantage function, which quantifies how much better the current action is compared to the baseline.
For clarity and space reasons, we omit detailed descriptions of the PPO algorithm, and instead present a summary of the training process in Fig 2, which outlines the key steps involved in training the PPO agent.
IV Experimental Results
IV-A Simulation Setup
The simulation environment is based on site-specific maps of the University of Southern California (USC) campus, covering an area of $512\times 512$ [m2]. This environment reflects dense urban features, including buildings and terrain, providing a realistic scenario to evaluate AutoBS.
Simulations are executed on a high-performance machine with a Tesla T4 GPU, 12GiB of RAM, and an Intel Xeon CPU @ 2.20GHz. Developed in Python with PyTorch for model training, the framework leverages PMNet for pathloss predictions.
Table I summarizes the key simulation parameters, including network configurations, environmental factors, and training setup.
Network Config.
Area $512\times 512$ [m2]
Carrier frequency $2.5$ [GHz]
Effective bandwidth $1$ [MHz]
Antenna Isotropic (vertical)
Input power 00 [dBm]
Noise figure $3$ [dB]
Coverage threshold $-90.015$ [dBm]
AutoBS Train Config.
DT generator PMNetusc [5]
DRL algorithm PPO [7]
Reward function Pathgain + Coverage ($\nu_{2}=0$)
Learning rate $1.0\times 10^{-5}$
Gamma $0.1$
SGD minibatch size $256$
# of samples for train (test) $1100$ ($50$)
# of episode $5$
Table I: Simulation Parameters.
Benchmark. To evaluate the performance of the proposed AutoBS framework, we compare it against two baseline methods: a heuristic-based deployment and an exhaustive search strategy. These baselines serve as the lower and upper bounds for performance comparison.
  • Heuristic: The heuristic deployment method randomly places BSs within the deployable area (i.e., $\mathcal{B}$), without incorporating learning or optimization. This simple approach serves as a performance lower bound, as it overlooks the impact of the site-specific information on the channel. It is also the assumption underlying stochastic-geometry system evaluations [8].
  • Exhaustive: The exhaustive search provides an upper bound for performance by systematically evaluating all possible BS placement configurations across the site-specific map. It guarantees a global optimal solution in terms of coverage or capacity based on the configuration criteria. However, the computational cost is exceedingly high, particularly for multi-BS deployments, as the complexity grows exponentially with the number of BSs and map size. Due to this limitation, the exhaustive search is restricted to $50$ samples in our baseline experiments. Additionally, in multi-BS deployments, the exhaustive method operates synchronously, unlike the asynchronous deployment strategy used by AutoBS. While an asynchronous exhaustive method could be defined to reduce computation time somewhat, it would still be computationally intense and would yield suboptimal results. Our focus remains on synchronous exhaustive deployment as it provides a true upper bound for comparison with AutoBS.
  • AutoBS (Proposed): The AutoBS leverages the PPO agent to make near-optimal BS deployment decisions in real time. It is trained using a fully connected neural network with four layers, each consisting of 128 nodes, to optimize both coverage and capacity. AutoBS achieves rapid deployment decisions at millisecond timescales.
(a) Building Map
(b) Heuristic
(c) Exhaustive
(d) AutoBS
Figure 3: Comparison of Single (Static) BS deployment in terms of coverage using SionnaRT. Light green areas indicate higher received signal strength. Fig. 3(a) shows a building map $\mathcal{S}$ used in each state $s_{t}$.
IV-B Performance Analysis
Scheme Coverage ($V$) Capacity ($C$)
Heuristic $43.37\%$ $1.892$
AutoBS $59.44\%$ $2.893$
Exhaustive for $V$ $64.23\%$ $2.981$
Exhaustive for $C$ $63.08\%$ $3.036$
Table II: Comparison results for Single (Static) BS deployment between exhaustive, random, and AutoBS deployment.
Comparison. Table II presents the performance of Heuristic, Exhaustive, and AutoBS deployments. AutoBS achieves significant improvements in both coverage and capacity compared to Heuristic deployment. Its performance closely approaches that of the Exhaustive method, which represents the global optimal solution, highlighting the effectiveness of policy-guided BS placement.
AutoBS focuses on coverage optimization through a reward function based on pathloss predictions from PMNet. Although capacity metrics are not explicitly included in the reward function (i.e., $\nu_{2}=0$), the observed capacity difference between AutoBS and Exhaustive search remains minimal—which will be further discussed in Sec. VI.
Coverage. Fig. 3 presents the coverage results for single BS deployment. The building map, representing the input state, includes key geographical features such as buildings and obstacles, which significantly affect BS placement and signal propagation.
In the Heuristic method, BSs are placed randomly across the deployable area, leading to inefficient coverage with noticeable gaps. This highlights the limitations of non-optimized deployments. The Exhaustive search method evaluates every possible BS placement option, achieving the best coverage by systematically considering all configurations, but at the price of much higher complexity.
The AutoBS method, trained with the PPO algorithm and PMNet for fast pathloss predictions, achieves coverage performance close to the exhaustive search but at a fraction of the computational cost.
It is important to note that AutoBS uses PMNet during training for fast and accurate pathloss map generation.  [^3] The plots are generated with SionnaRT purely because of its nicer graphical representation (the underlying numerical results are identical to those obtained with PMNet).
Convergence. Fig. 4 presents the convergence behavior for single BS deployment, comparing the AutoBS with Heuristic and Exhaustive methods (note that Heuristic and Exhaustive methods do not involve iterative learning). The AutoBS stabilizes between $50$ to $100$ training episodes, approaching the performance of the global optimal solution (i.e., Exhaustive), demonstrating its efficient learning and stable convergence.
Figure 4: Convergence behavior for Single (Static) BS deployment for Heuristic, Exhaustive, and AutoBS deployments over $200$ steps.
(a) Building map
(b) Heuristic
(c) Exhaustive
(d) AutoBS
Figure 5: Comparison results for Multi (Asynchronous) BS deployment in terms of coverage using SionnaRT. Light green areas indicate higher received signal strength. Fig. 5(a) shows a building map $\mathcal{S}$ used in each state $s_{t}$.
Multi-BS Deployment (Asynchronous). Table III and Fig. 5 present the results for asynchronous multi-BS deployment. The reward function is computed based only on the specific metrics (e.g., coverage, capacity, and/or pathgain) of each newly deployed BS. For example, when deploying the second BS, only the performance metrics for that BS are considered in the reward calculation.
For the multi-BS case, AutoBS performs an asynchronous deployment, i.e., adds one BS at a time, in a ”greedy” manner. This can be a practical requirement, e.g., when adding BSs to an existing infrastructure. In a greenfield deployment (like investigated here), the sequential approach will give worse results than a synchronous approach. Potential iterations to fine-tune already-placed BSs will be subject of our future research.
Scheme Coverage ($V$) Capacity ($C$)
Heuristic $63.95\%$ $2.966$
AutoBS $71.65\%$ $3.637$
Exhaustive for $V$ $81.06\%$ $3.995$
Exhaustive for $C$ $75.36\%$ $4.023$
Table III: Comparison results for Multi (Asynchronous) BS deployment between exhaustive, random, and AutoBS deployment.
While the overall findings align with those from the single (static) BS deployment, some notable insights emerge. One key observation is that deploying two BSs using the Heuristic method performs similarly or even worse than deploying a single BS with AutoBS or Exhaustive search in terms of coverage and capacity. This emphasizes the critical role of optimized deployment strategies. As shown in Table III for Heuristic and Table II for AutoBS and Exhaustive search, non-optimized deployments fall short even when more BSs are deployed.
These findings highlight two key aspects. First, AutoBS, despite using an asynchronous deployment strategy, achieves $90$% of the capacity of the global optimum (i.e., provided by the exhaustive method). Second, AutoBS’s ability to learn site-specific characteristics enables it to efficiently deploy BSs.
One BS Deploy. [sec] Two BS Deploy. [sec] Three BS Deploy. [sec]
Exhaustive 3.2245 ($\nicefrac{{10000}}{{10000}}$) 4062.1 ($\nicefrac{{1000000}}{{1000000}}$) 13449 ($\nicefrac{{1000000}}{{1000000}}$)
AutoBS 0.0145 ($\nicefrac{{45}}{{10000}}$) 0.1608 ($\nicefrac{{39}}{{1000000}}$) 0.2128 ($\nicefrac{{15}}{{1000000}}$)
Table IV: Time complexity analysis.
IV-C Complexity Analysis
Table IV highlights the time complexity analysis, emphasizing a key advantage of AutoBS: its ability to drastically reduce inference time, achieving millisecond-level decisions even in multi-BS deployments.
The efficiency gains are particularly notable in multi-BS deployments, where exhaustive search faces exponential complexity. AutoBS’s asynchronous deployment strategy enables it to scale seamlessly with growing network size, making it well-suited for large-scale networks and real-time optimization.
V Conclusion
This paper presents AutoBS, a RL-based framework for optimal BS deployment in 6G networks. By leveraging the PPO algorithm and integrating fast, accurate pathloss predictions from PMNet, AutoBS efficiently learns deployment strategies that adapt to site-specific channel dynamics, balancing coverage and capacity.
Our results show that AutoBS achieves up to $95\%$ of the capacity obtained by exhaustive search methods, while dramatically reducing inference time from hours to milliseconds. This makes AutoBS a practical solution for real-time deployments. With its ability to provide near-optimal performance and scale efficiently, AutoBS stands out as a powerful, automated solution tailored for the demands of large-scale 6G networks.
While AutoBS is designed for BS deployment in 6G networks, its framework can be easily adapted for other DT-based network optimization tasks, e.g., mobility management and link adaptation, by modifying the reward function and state information to match specific application goals. Furthermore, the flexible design of AutoBS makes it well-suited for additional applications, including Wi-Fi access point (AP) placement and O-RAN radio unit (RU) deployment.
VI Appendix: Reward Design
Reward Coverage ($V$) Capacity ($C$)
Capacity Only $57.642\%$ $2.862$
Coverage + Capacity $57.833\%$ $2.831$
Pathgain + Capacity $58.747\%$ $2.883$
Coverage Only $\textbf{59.537}\%$ $2.884$
Pathgain + Coverage $59.438\%$ 2.893
Table V: Reward design variations for single (static) BS deployment.
This section presents empirical findings from experiments with different reward designs in the AutoBS framework. The pathgain is defined as $P=\sum_{{i,j}\in\mathcal{R}}{P_{i,j}}$, and Table V summarizes coverage and capacity outcomes for each reward configuration. Although other transformations, such as logarithmic scaling (e.g., $\log{V}$ )), were tested, they are omitted here for brevity. These findings are empirical and sensitive to training hyperparameters (e.g., learning rate) and reward design parameters (e.g., $\nu$).
The results provide valuable insights. As expected, the Coverage Only reward achieves the highest coverage, aligning with its direct optimization objective (e.g., $\nu_{2},\nu_{3}=0$). Interestingly, in terms of capacity, the combination of Pathgain and Coverage (i.e., $\nu_{2}=0$) outperforms the Capacity Only reward. Notably, Coverage Only also surpasses Capacity Only in capacity, highlighting complex dynamics within the DRL training process. These findings suggest that site-specific channel characteristics are implicitly embedded in rewards derived from PMNet’s pathloss maps. While Capacity Only smooths variations through logarithmic scaling, Pathgain captures these fluctuations on a linear scale, providing more granular feedback for effective learning.
Throughout this work, the Pathgain + Coverage reward configuration is primarily used.
Acknowledgements: Part of this work was supported by Nokia Technologies and NSF-AoF grant 2133655. The authors gratefully acknowledge Arjun Balamwar and Yanqing Lu for their contributions to the framework design and simulations.